OpenStudy (anonymous):

The perimeter of a rectangle is the sum of all of its sides. Write a mathematical expression for the perimeter of a rectangle with length 4 and width w. Simplify the expression if possible

OpenStudy (diyadiya):

2(4+w)=8+2w

OpenStudy (diyadiya):

or 4+w + 4 +w =8 +2w

OpenStudy (anonymous):

Thanks.